Help me!
Rewrite the equation 4x-6y=-30 as a function of x.

Answer by funmath(2933)

This depends a little on your book and teacher.
Rewrite the equation 4x-6y=-30 as a function of x.
Solve the problem for y:
4x-4x-6y=-4x-30
-6y=-4x-30
-6y/-6=-4x/-6-30/-6
y=(2/3)x+5
Some teachers and books now have you replace y with f(x)
f(x)=(2/3)x+5
Happy Calculating!!!!